OpenStudy (hijacktrolo):

Charlie solved an equation, as shown below: Step 1: 5x = 30 Step 2: x = 30 – 5 Step 3: x = 25 Is Charlie's solution correct or incorrect? If the solution is incorrect, explain why it is incorrect and show the correct steps to solve the equation.

OpenStudy (hijacktrolo):

@DarrenMadx

OpenStudy (lolacole12):

hi

OpenStudy (hijacktrolo):

hiiiiiiiiii

OpenStudy (lolacole12):

Step 1: 5x = 30 Step 2: x = 30/5 Step 3: x = 6 B) the equation has 1 solution

OpenStudy (lolacole12):

charlie's answer is incorrect. in this case, you would divide 30/5 to isolate x. x=6. There is one answer

OpenStudy (hijacktrolo):

thanks

OpenStudy (lolacole12):

medal